# A unified framework for equation discovery and dynamic prediction of hysteretic systems

### v1

@@ 4.2 A complex structure with high-order nonlinear and fractional terms
To test the generalizability, the training and testing datasets are generated under different initial conditions. This experiment highlights the unified nature of the framework, as it remains applicable even when the underlying system exhibits higher-order nonlinearities and non-integer hysteretic exponents.
In this experiment, we consider two challenging cases as in Figure 1 and 2, namely Hysteresis Discovery and Full Equation Discovery, which differ in whether the dynamics motion equation ($f_{\theta}$) structure is assumed to be known (the structure of the hysteretic link equation ($g_{\phi}$) is assumed to be unknown in both cases). This comparison enables us to examine the impact of prior structural knowledge on equation discovery for complex hysteretic systems. In this experiment, five different initial conditions are used for training, while an additional unseen initial condition is used for testing to evaluate the generalizability of the proposed framework. The external excitation is defined as $u(t)=\sin(2t)$, representing a smooth periodic input suitable for dynamic identification. The sampling frequency is set to $f_{s}=100~\text{Hz}$, and the total simulation duration is $T=6s$. Such a setup allows for a comprehensive assessment of the model’s robustness under varying initial states while maintaining consistent excitation characteristics.
@@
Figure 7 shows the hysteresis loops learned by the solver-based internal variable learning module. Both configurations can capture qualitative hysteretic behavior, yet the Hysteresis Discovery model clearly achieves better accuracy in reconstructing the loop dynamics. Table 3 lists the explicit forms of the discovered equations, and Table 4 reports the NRMSE results. Figures 8 and 9 compare the displacement, velocity, and acceleration responses obtained by solving the discovered equations, compared with the corresponding ground-truth responses in the training and testing datasets, respectively.
@@
From these results, it is evident that the proposed framework consistently outperforms SINDy, achieving lower errors and more faithful recovery of the governing equations. For SINDy, the pre-defined library was designed to be as consistent as possible with our method: the dynamics motion equation is assumed to be a linear combination of $x$, $\dot{x}$, $z$, and $u$, while the hysteretic link equation is constrained to contain absolute values and their integer powers (from 1 to 5), to provide SINDy with sufficient expressive capacity for capturing nonlinear hysteretic behaviors. However, unlike SR, SINDy cannot flexibly capture fractional exponents.
In particular, both Hysteresis Discovery and Full Equation Discovery can successfully recover expressions with fractional exponents in Eq.(12), which is generally beyond the capability of library-based approaches such as SINDy. By contrast, the Full Equation Discovery setting—where both the dynamics motion equation and the hysteretic link equation are unknown—yields slightly lower accuracy than Hysteresis Discovery. In practice, it is suggested that incorporating some prior knowledge helps improve accuracy and equation discovery performance.
Table 3: Discovered equations results of complex structure data ($t$ is omitted for brevity). The true equations are $1.0\ddot{x}+0.8\dot{x}+0.5x^{3}+1.0z=u$ and $\dot{z}=4.0\dot{x}-5.0|\dot{x}||z|^{0.5}z+4.0\dot{x}|z|^{1.5}$. SINDy fails to recover the fractional order terms and instead introduces spurious integer order terms to compensate.
Type and method
Discovered equations
Hysteresis Discovery
$0.9388\ddot{x}+0.7934\dot{x}+0.5366x^{3}+1.0320z=u$
$\dot{z}=4.0633\dot{x}-4.8310|\dot{x}||z|^{0.5486}z+3.8942\dot{x}|z|^{1.5486}$
Hysteresis Discovery
$\ddot{x}+0.8445\dot{x}+0.5679x^{3}+1.0992z=1.0649u$
$\dot{z}=-6.80922\,|\dot{x}|\,|z|z+5.00018\,|\dot{x}|\,|z|^{2}z+4.98652\,\dot{x}\,|z|^{2}+3.73699\,\dot{x}$
Full Equation
Discovery (Ours)
$\ddot{x}+0.7689\dot{x}+0.4573x^{3}+1.0423z=0.9503u$
$\dot{z}=4.0543\dot{x}-5.1348|\dot{x}||z|^{0.6225}z+3.8224\dot{x}|z|^{1.6225}$
Full Equation
Discovery (SINDy)
$\ddot{x}+0.7490\dot{x}+0.5833x^{3}+1.2145z=0.8992u$
$\dot{z}=-12.704|\dot{x}||z|z+6.118|\dot{x}||z|^{2}z+2.941\dot{x}|z|^{2}-0.873\dot{x}$
@@
Type Responses
Training
NRMSE
Testing
NRMSE
Training
NRMSE
Testing
NRMSE
Hysteresis
discovery
Displacement $x$ 4.21% 3.84% 25.42% 36.87%
Velocity $\dot{x}$ 3.91% 4.01% 24.79% 39.50%
Acceleration $\ddot{x}$ 3.86% 3.88% 25.70% 40.67%
Full equation
discovery
Displacement $x$ 5.28% 4.57% 25.73% 37.76%
Velocity $\dot{x}$ 4.56% 4.25% 26.56% 32.43%
Acceleration $\ddot{x}$ 4.86% 4.31% 23.00% 36.15%
@@ 4.3 An experimental SDOF yielding structure
Figure 10: Shaking table test of a braced 3-story structure (equivalent to a SDOF system), with a negative stiffness device (NSD) installed on the first floor [lai2019sparse].
@@
In this experiment, a shear-type SDOF yielding structure excited by shake-table motions is implemented to investigate the proposed method. The experimental setup of the braced 3-story structure (equivalent to a SDOF system), with and without a negative stiffness device (NSD) [sarlis2013negative, pasala2013adaptive] is illustrated in Figure 10 [lai2019sparse, pasala2013seismic, Attary17022015], where the NSD is installed on the first floor to alter hysteretic behaviors.
@@
We adopt the governing equation structure $m\ddot{x}(t)+c\dot{x}(t)+kx(t)+\alpha z(t)=-mu(t)$, where $x(t)$ is the displacement, $\dot{x}(t)$ and $\ddot{x}(t)$ are its velocity and acceleration, $u(t)$ is the measured base acceleration, and $z(t)$ is the hysteretic internal variable. This setup belongs to the hysteresis discovery type, in which the structure of the equation of motion is predefined (which needs to identify unknown system parameters), and the nonlinear hysteretic link equation is assumed to be unknown.
@@
We reuse the real shake table experiment from [lai2019sparse], and study the structure in two scenarios: (i) without an NSD and (ii) with an NSD installed on the first floor. Following the setup in [lai2019sparse], we use the measured displacement $x(t)$, the computed velocity $\dot{x}(t)$ (via differentiation and smoothing), the measured acceleration $\ddot{x}(t)$, and the shake-table excitation $u(t)$ as inputs to our framework. For the without NSD case, a type-I 60% scaled Pacoima earthquake is used for training, and a reversed type-II 60% Pacoima record for testing. For the with NSD case, the training signal concatenates 43% and 65% scaled Pacoima segments, while the testing input is a 55% scaled Sylmar record (see Figs. 23–25 of [lai2019sparse] for details).
For the without NSD case, the hysteresis loop in the $x˙\dot{x}$ plane is presented in Figure 11(a). The discovered evolution law reduces to rate-dominated form $\dot{z}(t)=2.1847\dot{x}(t)$, together with the dynamics motion equation $22.2798\ddot{x}(t)+18.4829\dot{x}(t)+1571.7874x(t)+1572.5618z(t)=-22.2798u(t)$, which in nature is a linear system. This reflects that our proposed framework is capable of automatically discovering a linear equation if the system comes without evident nonlinearity or hysteresis. Figures 12 and 13 show that the model closely matches the training and testing responses in terms of displacement, velocity, and acceleration. Table 5 provides quantitative results. The small NRMSEs confirm that the framework achieves accurate and robust identification when the hysteresis is mild and rate-dominated.
For the with NSD case, the hysteresis loops become much wider and exhibit obvious pinching and asymmetry (Figure 11(b)). Accordingly, the learned $\dot{z}(t)$ includes higher-order and mixed-magnitude terms of $x(t)$, $\dot{x}(t)$, and $z(t)$. The predicted hysteretic link equation is $\dot{z}(t)=4.8586\dot{x}(t)+1.6412x^{2}|\dot{x}(t)|+1.8160\dot{x}(t)|\dot{x}(t)|z(t)-1.8174\dot{x}(t)|z(t)|+1.7648z(t)|z(t)|$ with the dynamics motion equation $2.2798\ddot{x}(t)+44.7281\dot{x}(t)+2099.6526x(t)+2108.1516z(t) =22$.2798u(t)=-22.2798u(t). As shown in Figures 14 and 15, the training responses are well captured, but testing displacement exhibits larger deviations. Table 5 confirms this: the displacement error rises to 14.11% in the testing set, while velocity and acceleration remain within 6%. This indicates that the NSD introduces stronger nonlinear hysteresis, which requires a richer $\dot{z}$ structure and leads to a larger generalization gap. In other words, the model still succeeds in capturing the main nonlinear patterns, but the predictive accuracy is more sensitive to unseen excitations.
From all these results, it can be concluded that the proposed framework is capable of both and equation discovery and dynamic prediction in real experimental data. For the relatively simple without NSD case, it achieves concise and highly accurate identification, while for the more challenging with NSD case, it still provides interpretable governing equations and meaningful forward predictions, albeit with a higher generalization error. In future work, these limitations may be further addressed by extending the framework toward more complex hysteretic systems and improving its generalization to a wider range of nonlinear dynamical problems.
@@
Scenarios Responses
Training
NRMSE
Testing
NRMSE
Training
NRMSE
(Sparse [lai2019sparse])
Testing
NRMSE
(Sparse [lai2019sparse])
Without NSD
Displacement $x$ 3.23% 2.99% 10.73% 2.90%

### v2

@@ 4.2 A complex structure with high-order nonlinear and fractional terms
This experiment highlights the unified nature of the framework, as it remains applicable even when the underlying system exhibits higher-order nonlinearities and non-integer hysteretic exponents.
Before presenting the main experimental results, we first examine the sensitivity of the learned hysteretic variable to initialization in the most challenging Full Equation Discovery setting in this experiment. Specifically, we consider a fixed training case with only one initial condition and one excitation, while initializing the trainable hysteretic variable $\mathbf{z}$ through Step 1 in three different strategies: (a) Ours: initialization described in Section 3.2, (b) Zero: $z(t)=0$, (c) Sine: $z(t)=\sin(2\pi t)$. Figure 7 shows the xxzz hysteresis loops for these three strategies at initialization and after convergence. Figure 8 shows the Step 1 predicted results, it can be seen that all three initialization strategies can accurately predict $x$, $\dot{x}$, and $\ddot{x}$.
However, the strategies Zero and Sine struggle to recover the correct hysteresis loops, indicating that accurate response prediction does not necessarily imply accurate learning of the hysteretic variable. Although the initial hysteresis loop delivered by our initialization strategy differs from the ground truth, it captures hysteretic features compared with the other two strategies. After training, the final optimized results become much closer to the ground truth.
This result suggests that, when the training data are not sufficiently informative, strict uniqueness of the learned hysteretic variable cannot be asserted. Nevertheless, the proposed solver-based learning scheme remains a practically effective initialization strategy. This example represents a relatively extreme case, as only a single training trajectory is used. It also highlights the importance of sufficient diversity in the training data, which is examined in this experiment by considering multiple initial conditions.
Figure 7: xxzz hysteresis loops under the same single initial condition and excitation in the Full Equation Discovery setting of complex structure data. Three different strategies: (a) Ours: initialization described in Section 3.2. (b) Zero: $z(t)=0$. (c) Sine: $z(t)=\sin(2\pi t)$.
Figure 8: Predicted response results in Step 1 of complex structure data. (a) Displacement results. (b) Velocity results. (c) Acceleration results.
The external excitation is defined as $u(t)=\sin(2t)$, representing a smooth periodic input suitable for a dynamic system. The sampling frequency is set to $f_{s}=100~\text{Hz}$, and the total simulation duration is $T=6\text{s}$. Such a setup allows for a comprehensive assessment of the model’s robustness under varying initial states while maintaining consistent excitation characteristics.
To compare the proposed framework with a representative sparse regression pipeline, we use SINDy [6], which discovers governing equations by performing sparse regression over a pre-defined candidate function library. In addition to the vanilla SINDy, we include two advanced SINDy methods to provide a fairer comparison under strong nonlinearity and noisy measurements. Specifically, we consider parallel implicit SINDy (SINDy-PI) [21], which identifies implicitly defined governing relations and leverages multiple trajectories to improve identifiability, and Weak SINDy [30], which performs identification in an integral (weak) form to reduce sensitivity to derivative noise. For all SINDy-based methods, we use a candidate library design that is made as consistent as possible with our method, and we tune the key sparsity-related hyperparameters on the training set to avoid biasing the comparison. Specifically, the dynamic motion equation is assumed to be a linear combination of $x$, $\dot{x}$, $z$, and $u$, whereas the hysteretic link equation is constrained to include absolute-value terms and their integer powers from 1 to 5, thereby providing SINDy with sufficient expressive capacity to capture nonlinear hysteretic behaviors.
@@
Figure 9 shows the hysteresis loops learned by the proposed framework. Both configurations can capture qualitative hysteretic behavior, yet the Hysteresis Discovery model clearly achieves better accuracy in reconstructing the loop dynamics. Table 3 lists the explicit forms of the discovered equations, and Table 4 reports the NRMSE results
Figures 10 and 11 compare the displacement, velocity, and acceleration responses obtained by solving the discovered equations, compared with the corresponding ground truth responses in the training and testing datasets, respectively.
@@
From these results, it is evident that the proposed framework consistently outperforms SINDy-based methods, achieving lower errors and more faithful recovery of the governing equations. unlike SR, SINDy cannot flexibly capture fractional exponents. both Hysteresis Discovery and Full Equation Discovery can successfully recover expressions with fractional exponents, which is generally beyond the capability of library-based approaches such as SINDy. By contrast, the Full Equation Discovery setting—where both the dynamics motion equation and the hysteretic link equation are unknown—yields slightly lower accuracy than Hysteresis Discovery. In practice, it is suggested that incorporating some prior knowledge helps improve accuracy and equation discovery performance.
Table 3: Discovered equations results of complex structure data ($t$ is omitted for brevity). SINDy fails to recover the fractional-order terms and instead introduces spurious integer-order terms to compensate.
Type and method Discovered equations
Hysteresis Discovery (Ours) $0.9388\ddot{x}+0.7934\dot{x}+0.5366x^{3}+1.0320z=u$ $\dot{z}=4.0633\dot{x}-4.8310|\dot{x}|||^{0.5486}z+3.8942\dot{x}|z|^{1.5486}$
Hysteresis Discovery (SINDy) $\ddot{x}+0.8445\dot{x}+0.5679x^{3}+1.0992z=1.0649u$ $\dot{z}=-6.80922\,|\dot{x}|\,|z|z+5.00018\,|\dot{x}|\,|z|^{2}z+4.98652\,\dot{x}\,|z|^{2}+3.73699\,\dot{x}$
Full Equation Discovery (Ours) $\ddot{x}+0.7689\dot{x}+0.4573x^{3}+1.0423z=0.9503u$ $\dot{z}=4.0543\dot{x}-5.1348|\dot{x}|||^{0.6225}z+3.8224\dot{x}|z|^{1.6225}$
Full Equation Discovery (SINDy) $\ddot{x}+0.7490\dot{x}+0.5833x^{3}+1.2145z=0.8992u$ $\dot{z}=-12.704|\dot{x}||z|z+6.118|\dot{x}||z|^{2}z+2.941\dot{x}|z|^{2}-0.873\dot{x}$
@@
Type Responses Method Training NRMSE Testing NRMSE
Hysteresis Discovery Displacement $x$ Ours 4.21% 3.84%
SINDy 25.42% 36.87%
Velocity $\dot{x}$ Ours 3.91% 4.01%
SINDy 24.79% 39.50%
Acceleration $\ddot{x}$ Ours 3.86% 3.88%
SINDy 25.70% 40.67%
Full Equation Discovery Displacement $x$ Ours 5.28% 4.57%
SINDy 25.73% 37.76%
Velocity $\dot{x}$ Ours 4.56% 4.25%
SINDy 26.56% 32.43%
Acceleration $\ddot{x}$ Ours 4.86% 4.31%
SINDy 23.00% 36.15%
@@ 4.3 An experimental SDOF yielding structure
Figure 12: Shaking table test of a braced 3-story structure (equivalent to a SDOF system), with a negative stiffness device (NSD) installed on the first floor [23].
@@
In this experiment, a shear-type SDOF yielding structure excited by shake-table ground motions is implemented to investigate the proposed framework method. The experimental setup of the braced 3-story structure (equivalent to a SDOF system), with and without a negative stiffness device (NSD) [39, 35] is illustrated in Figure 12 [23, 34, 2], where the NSD is installed on the first floor to alter hysteretic behaviors.
@@
We adopt the governing equation structure $m\ddot{x}(t)+c\dot{x}(t)+kx(t)+\alpha z(t)=-mu(t)$, where $x(t)$ is the displacement, $\dot{x}(t)$ and $\ddot{x}(t)$ are its velocity and acceleration, $u(t)$ is the measured base acceleration, and $z(t)$ is the hysteretic variable. This setup belongs to the hysteresis discovery type, in which the structure of the equation of motion is predefined (which needs to identify unknown system parameters), and the nonlinear hysteretic link equation is assumed to be unknown.
@@
We reuse the real shake table experiment from [23], and study the structure in two scenarios: (i) without an NSD and (ii) with an NSD installed on the first floor. Following the setup in [23], we use the measured displacement $x(t)$, the computed velocity $\dot{x}(t)$ (via differentiation and smoothing), the measured acceleration $\ddot{x}(t)$, and the shake-table excitation $u(t)$ as inputs to our framework. For the without NSD case, a type-I 60% scaled Pacoima earthquake is used for training, and a reversed type-II 60% Pacoima record for testing. For the with NSD case, the training signal concatenates 43% and 65% scaled Pacoima segments, while the testing input is a 55% scaled Sylmar record (see Figs 23–25 of [23] for details).
@@
For the without NSD case, the hysteresis loop in the $x˙\dot{x}$ plane is presented in Figure 13 (a). The discovered evolution law reduces to a rate-dominated form $\dot{z}(t)=2.1847\dot{x}(t)$, together with the motion equation $22.2798\ddot{x}(t)+18.4829\dot{x}(t)+1571.7874x(t)+1572.5618z(t)=-22.2798u(t)$, which in nature is a linear system. This reflects that our proposed framework is capable of automatically discovering a linear equation if the system comes without evident nonlinearity or hysteresis. Figures 14 and 15 show that the model closely matches the training and testing responses in terms of displacement, velocity, and acceleration. Table 5 provides quantitative results. The small NRMSEs confirm that the framework achieves accurate and robust identification when the hysteresis is mild and rate-dominated.
For the with NSD case, the hysteresis loops become much wider and exhibit obvious pinching and asymmetry (Figure 13 (b)). Accordingly, the learned $\dot{z}(t)$ includes higher-order and mixed-magnitude terms of $x(t)$, $\dot{x}(t)$, and $z(t)$. The predicted hysteretic link equation is $\dot{z}(t)=4.8586\dot{x}(t)+1.6412x^{2}|\dot{x}(t)|+1.8160\dot{x}(t)|\dot{x}(t)|z(t)-1.8174\dot{x}(t)|z(t)|+1.7648z(t)|z(t)|$ with the dynamics motion equation $2.2798\ddot{x}(t)+44.7281\dot{x}(t)+2099.6526x(t)+2108.1516z(t) =22$.2798u(t)=-22.2798u(t). As shown in Figures 16 and 17, the training responses are well captured, but testing displacement exhibits larger deviations. Table 5 confirms this: the displacement error rises to 14.11% in the testing set, while velocity and acceleration remain within 6%. This indicates that the NSD introduces stronger nonlinear hysteresis, which requires a richer $\dot{z}$ structure and leads to a larger generalization gap. In other words, the model still succeeds in capturing the main nonlinear patterns, but the predictive accuracy is more sensitive to unseen excitations.
From all these results, it can be concluded that the proposed framework is capable of both equation discovery and dynamic prediction in real experimental data. For the relatively simple without NSD case, it achieves concise and highly accurate identification, while for the more challenging with NSD case, it still provides interpretable governing equations and meaningful forward predictions, albeit with a higher generalization error. In future work, these limitations may be further addressed by extending the framework toward more complex hysteretic systems and improving its generalization to a wider range of nonlinear dynamical problems.
@@
Scenarios Responses Training NRMSE (Ours) Testing NRMSE (Ours) Training NRMSE (Sparse [23]) Testing NRMSE (Sparse [23])
Without NSD Displacement $x$ 3.23% 2.99% 10.73% 2.90%
@@
4.4 Extension to MDOF systems
The proposed framework for equation discovery and dynamic prediction is not restricted to SDOF hysteretic systems. For an MDOF system [26, 27], we promote the vector states to the matrix and represent the coupled dynamics ($t$ is omitted for brevity in this section) with $r$-DOF and $q$ hysteretic variable(s). Therefore, the equations can be written as:
$\left\{\begin{aligned} &\boldsymbol{M}\ddot{\boldsymbol{x}}+\boldsymbol{C}\dot{\boldsymbol{x}}+\boldsymbol{K}\boldsymbol{x}+\boldsymbol{H}\boldsymbol{z}=\boldsymbol{B}\boldsymbol{u},\\ &\dot{\boldsymbol{z}}=\boldsymbol{g}_{\phi}\!\left(\dot{\boldsymbol{x}},\boldsymbol{z}\right),\end{aligned}\right$. (17)
where $\boldsymbol{x}\in\mathbb{R}^{r}$ denotes the displacement vector (with $r$-DOF), and $\boldsymbol{z}\in\mathbb{R}^{q}$ collects the hysteretic variables associated with localized hysteretic elements (e.g., inter-story links, yielding dampers, or absorber components). The matrices $\boldsymbol{M},\boldsymbol{C},\boldsymbol{K}\in\mathbb{R}^{r\times r}$ are the mass, damping, and stiffness matrices, respectively. The coupling matrix $\boldsymbol{H}\in\mathbb{R}^{r\times q}$ maps the hysteretic variables to generalized restoring forces, while $\boldsymbol{B}\in\mathbb{R}^{r\times p}$ distributes the external input $\boldsymbol{u}\in\mathbb{R}^{p}$ to the structural DOFs. The (possibly nonlinear) function $\boldsymbol{g}_{\phi}$ governs the hysteretic variable evolution and can encode rate dependence and memory effects.
To illustrate the above formulation, we consider a simple yet representative case with $r=2,q=1$ and $p=1$, corresponding to a 2-DOF model equipped with a single hysteretic energy-dissipating link. Let:
$\boldsymbol{x}=\begin{bmatrix}x_{1}\\ x_{2}\end{bmatrix},\qquad z\in\mathbb{R},$ (18)
where $x_{1}$ and $x_{2}$ are the displacements of the 2-DOFs, and $z$ represents the hysteretic variable. The coupled dynamics in Eq.(17) can be reduce to:
$\left\{\begin{aligned} &\boldsymbol{M}\ddot{\boldsymbol{x}}+\boldsymbol{C}\dot{\boldsymbol{x}}+\boldsymbol{K}\boldsymbol{x}+\boldsymbol{H}z=\boldsymbol{B}u,\\ &\dot{z}=g_{\phi}\!\left(\dot{x}_{1},\dot{x}_{2},z\right).\end{aligned}\right$. (19)
As a concrete numerical example, we choose a harmonic excitation $u(t)=1000\sin\!\left(2\pi\times 1.5t\right)$ and $\textbf{B}=[1,1]^{T}$, specify:
$\boldsymbol{M}=\begin{bmatrix}m_{1}&0\\ 0&m_{2}\end{bmatrix}=\begin{bmatrix}100&0\\ 0&80\end{bmatrix},\quad\boldsymbol{C}=\begin{bmatrix}c_{11}&c_{12}\\ c_{21}&c_{22}\end{bmatrix}=\begin{bmatrix}30&-5\\ -5&20\end{bmatrix}$, (20)
$\boldsymbol{K}=\begin{bmatrix}k_{11}&k_{12}\\ k_{21}&k_{22}\end{bmatrix}=\begin{bmatrix}3500&-1500\\ -1500&1500\end{bmatrix},\quad\boldsymbol{H}=\begin{bmatrix}h_{1}\\ h_{2}\end{bmatrix}=\begin{bmatrix}1\\ -1\end{bmatrix}$. (21)
With this choice, the hysteretic link exerts equal and opposite restoring forces on the 2-DOFs. For the hysteretic variable evolution, we adopt a simple rate-dependent form driven by the inter-story relative velocity $\dot{x}_{2}-\dot{x}_{1}$:
$\dot{z}=a(\dot{x}_{2}-\dot{x}_{1})-b\left|\dot{x}_{2}-\dot{x}_{1}\right|\,z,\qquad a=50,\;\;b=2.$ (22)
For completeness, Eq.(19) can be written component-wise as:
$\left\{\begin{aligned} &100\,\ddot{x}_{1}+30\,\dot{x}_{1}-5\,\dot{x}_{2}+3500\,x_{1}-1500\,x_{2}+z=u,\\ &80\,\ddot{x}_{2}-5\,\dot{x}_{1}+20\,\dot{x}_{2}-1500\,x_{1}+1500\,x_{2}-z=u,\\ &\dot{z}=50(\dot{x}_{2}-\dot{x}_{1})-2\left|\dot{x}_{2}-\dot{x}_{1}\right|\,z,\end{aligned}\right$. (23)
which represent the dynamic motion equation $f_{\theta_{1}},f_{\theta_{2}}$ and the hysteretic link equation $g_{\phi}$, respectively.
The above derivations illustrate that the adopted state-space formulations extend naturally to MDOF settings by treating both structural states and hysteretic variables in a coupled form. Therefore, the formulation remains structurally identical to the SDOF case, with the equation discovery task now applied to Eq.(23). In what follows, we present the Hysteresis Discovery case in Figure 1, where the measurements are noise-free and $\mathbf{x}$, $\dot{\mathbf{x}}$, and $\ddot{\mathbf{x}}$ are all known, and the other settings are similar. Combining the procedure in the proposed framework, the unknowns include the entries of $\mathbf{M}$, $\mathbf{C}$, $\mathbf{K}$, and $\mathbf{H}$, as well as the entire functional form of the hysteretic-variable evolution in Eq.(22). Using the strategy of reformulation introduced in Section 3, the MDOF formulation can be written explicitly in a first-order state-space form:
$\frac{d}{dt}\begin{bmatrix}x_{1}\\ x_{2}\\ \dot{x}_{1}\\ \dot{x}_{2}\end{bmatrix}=\begin{bmatrix}\dot{x}_{1}\\ \dot{x}_{2}\\ f_{\theta_{1}}\\ f_{\theta_{2}}\end{bmatrix}$, (24)
with $z$ being trainable.
The discovered governing equations and NRMSE results can be seen in Table LABEL:tab:mdof_eq and LABEL:tab:mdof_nrmse. Figure 18 shows the training and testing results of the discovered MDOF system for both displacement and velocity responses. These results indicate that the proposed framework remains effective when extended from SDOF to MDOF systems, yielding reasonably accurate equation discovery and dynamic predictions.
We acknowledge that, as in many MDOF settings, the problem becomes increasingly complex as the number of DOF grows. As $r$ and $q$ increase, both computational cost and identifiability challenges can become more pronounced, especially when hysteresis is present in multiple or adjacent DOFs due to structural coupling. Nevertheless, based on above observations, the proposed framework remains directly applicable and can handle this increased complexity.
[…]